OpenStudy (anonymous):

What is the value for x in the proportion ? x over x+10 =7 over 12 Numerical Answers Expected!

jimthompson5910 (jim_thompson5910):

x/(x+10) = 7/12 12x = 7(x+10) 12x=7x+70 12x-7x=70 5x=70 x=(70)/(5) x = 14